If the volume of a sphere is 500pi/3 meters cubed, what is the diameter of the sphere?
Answer by jim_thompson5910(35256)


V = (4/3)*pi*r^3

500pi/3 = (4/3)*pi*r^3

500/3 = (4/3)*r^3 ... Divide both sides by pi.

500 = 4r^3 ... Multiply both sides by 3.

500/4 = r^3

125 = r^3

r^3 = 125

r = cube root of 125

r = 5

The radius is 5 meters. So the diameter is 10 meters.